### v1

# Evaluating AI for Finance: Is AI Credible at Assessing Investment Risk Appetite?

Abstract
We evaluate the credibility of leading AI models in assessing investment risk appetite. Our analysis spans proprietary (GPT-4, Claude 3.7, Gemini 1.5) and open-weight models (LLaMA 3.1/3.3, DeepSeek-V3, Mistral-small), using 1,720 user profiles constructed with 16 risk-relevant features across 10 countries and both genders. We observe significant variance across models in score distributions and demographic sensitivity. For example, GPT-4o assigns higher risk scores to Nigerian and Indonesian profiles, while LLaMA and DeepSeek show opposite gender tendencies in risk classification. While some models (e.g., GPT-4o, LLaMA 3.1) align closely with expected scores in low- and mid-risk ranges, none maintain consistent performance across regions and demographics. Our findings highlight the need for rigorous, standardized evaluations of AI systems in regulated financial contexts to prevent bias, opacity, and inconsistency in real-world deployment.
Evaluating AI for Finance: Is AI Credible at Assessing Investment Risk Appetite?
Divij Chawlathanks: Equal contribution,, Ashita Bhutada*, Do Duc Anh, Abhinav Raghunathan, Vinod SP, Cathy Guo, Dar Win Liew, Prannaya Gupta, Rishabh Bhardwajthanks: Lead contributors, email: rishabh@walled.ai, Rajat Bhardwaj, Soujanya Poria Walled AI Labs
1 Introduction
Artificial intelligence (AI), particularly generative AI powered by large language models (LLMs), is rapidly reshaping multiple industries. These AI technologies assist with a variety of complex tasks—from drafting emails and writing code to synthesizing research—activities that traditionally required significant time and domain expertise. This automation promises to improve efficiency and productivity dramatically.
Recent industry reports emphasize the scale and pace of this transformation. Gartner Gartner (2023) projects that by 2026, over 80% of enterprises will adopt generative AI tools or deploy AI-powered applications. Similarly, IDC IDC (2024) reports that 92% of organizations using AI have seen significant productivity gains, with average returns on investment reaching 3.7 times their initial cost, with some experiencing returns up to tenfold. While these developments underscore AI’s transformative potential, they also introduce new challenges—particularly in high-stakes, regulated domains such as finance.
Why Focus on AI in Finance?
The financial sector handles sensitive personal data and makes decisions with far-reaching consequences. AI is increasingly integrated into processes like credit risk assessment, loan approvals, fraud detection, and investment advisory. However, AI models are not flawless. Inaccurate or biased predictions from these systems can result in serious harms: unfair loan denials, misallocation of capital, discrimination against certain demographic groups, and breaches of regulatory compliance.
For example, an AI model that inaccurately estimates an individual’s willingness or ability to take investment risks might recommend unsuitable financial products, potentially causing financial harm. Moreover, if the AI exhibits bias—assigning systematically higher or lower risk tolerance based on gender, ethnicity, or nationality—it can exacerbate existing financial inequalities and damage trust in AI systems. Regulators worldwide are increasingly scrutinizing AI tools to ensure fairness, transparency, and accountability.
Defining Key Concepts: Risk Appetite and Credibility
Before we delve further, it is important to define two core concepts central to this study:
  • Investment Risk Appetite (or Risk Tolerance): This refers to an investor’s willingness and capacity to endure financial losses or volatility in pursuit of potential returns. It is a critical factor guiding investment decisions and portfolio allocation. Accurate risk appetite assessment enables financial advisors to recommend products aligned with individual preferences and needs.
  • Credibility of AI Models: In this context, credibility refers to the trustworthiness of an AI model’s risk appetite predictions. We consider two dimensions:
    1. 1.
      Correctness: The extent to which the AI’s risk scores align with expected or “ideal” risk profiles, given known financial and demographic attributes information.
    2. 2.
      Consistency: The stability of AI predictions across user characteristics, such as gender and nationality. Consistency ensures the model behaves fairly and does not exhibit demographic-based variation for otherwise similar profiles.
Research Questions and Stakes
This study seeks to rigorously evaluate the credibility of current AI models in assessing investment risk appetite by addressing three core questions:
  1. 1.
    Accuracy: Do AI models produce risk appetite scores that correctly reflect users’ financial situations and preferences?
  2. 2.
    Consistency: Are these predictions stable across users from different regions and demographic groups?
  3. 3.
    Bias: Do AI systems exhibit systematic bias that favor or disadvantage users based on gender or country of origin?
Addressing these questions is crucial. Financial institutions and regulators must trust that AI systems operate fairly and reliably. Failure to detect or mitigate inaccuracies and biases could cause harmful outcomes, such as unfair treatment of clients, financial losses, reputational harm, and regulatory sanctions. Given the sensitive and high-stakes nature of finance, rigorous evaluation is not optional—it is essential.
Our Approach
To evaluate AI models credibility, we designed a comprehensive experimental framework:
  • We created a dataset of 1,720 synthetic user profiles. Each profile includes 16 carefully selected features related to financial status, investment goals, and other risk-relevant characteristics. These features are grouped into categories such as Financial Stability, Income, Investment Objectives, and more (see Figure 1).
  • The profiles capture a diverse population across 10 countries and maintain balanced gender representation. This diversity allows us to test AI models on a wide range of realistic financial scenarios and demographic groups.
  • We evaluated several leading AI models, including proprietary systems like GPT-4, Claude 3.7, and Gemini 1.5, as well as open-weight models like LLaMA 3.1/3.3, DeepSeek-V3, and Mistral-small.
  • Each model was prompted to generate an investment risk appetite score for every user profile. We then analyzed the results to assess how well each model performed in terms of correctness, consistency, and bias.
Key Findings
Our analysis uncovers significant variability and limitations in current AI systems:
  • Models such as GPT-4o demonstrate strong alignment with expected risk tolerance scores for low- and mid-risk profiles.
  • However, some models show demographic biases. For example, GPT-4o tends to assign higher risk scores to Nigerian and Indonesian profiles, while open-weight models like LLaMA and DeepSeek display inconsistent gender-based scoring trends.
  • No AI system consistently produces unbiased, accurate risk scores across all demographic groups and countries.
These findings highlight the pressing need for:
  • Standardized evaluation protocols to rigorously benchmark AI systems on fairness and accuracy.
  • Improved AI training and calibration methods to reduce bias and enhance reliability.
  • Transparent reporting mechanisms to foster trust among users and regulators.
Paper Structure
The remainder of this paper is organized as follows:
  • Section 2 describes the selection and significance of risk-relevant features in our profiles.
  • Section 3 explains our methodology for scoring and evaluating AI predictions.
  • Section 4 outlines data generation and profile construction.
  • Section 5 details the experimental setup and model prompting procedures.
  • Section 6 presents the results, focusing on correctness, consistency, demographic biases, and implications for financial AI deployment.
2 What Features Are Important for Risk Profiling?
To benchmark AI models on investment risk assessment, we first identify the most relevant user features. Our selection draws from regulatory standards, academic research, and industry practices. Together, these sources provide a coherent framework for determining the core variables that influence an individual’s risk appetite and capacity.
Regulatory Frameworks.
Global financial regulators consistently emphasize the need to evaluate an individual’s financial circumstances, objectives, and knowledge when assessing risk tolerance.
The United States’ Financial Industry Regulatory Authority (FINRA) highlights the importance of income, assets, liabilities, and investment objectives when evaluating investor suitability FINRA (2012). The United Kingdom’s Financial Conduct Authority (FCA) recommends considering financial position, investment experience, and risk preferences to ensure that investment advice is appropriate FCA (2018). Similarly, the European Securities and Markets Authority (ESMA) emphasizes suitability assessments that account for both the financial situation and investment goals of the client ESMA (2023).
Japan’s Financial Services Agency (FSA) calls for evaluation of an investor’s financial stability, loss-bearing capacity, and investment literacy FSA (2022). The Monetary Authority of Singapore (MAS) requires financial advisors to assess income, assets, liabilities, and investment objectives as core components of risk profiling MAS (2023).
Despite geographic variation, these regulators converge on a common set of variables: personal finances, investment goals, and user knowledge or experience. These factors serve as the regulatory foundation for any credible risk assessment.
Figure 1: Factors determining investment risk profile.
Research and Industry Frameworks.
Beyond regulation, we examined academic literature and best practices from global investment firms with operations spanning over ten countries and more than fifty years of service.
Personal Financial Stability—comprising stable income and manageable liabilities—is a primary determinant of risk capacity Grable and Lytton (1999). An individual with stable income and low debt can typically tolerate higher risk, as they are better insulated against short-term market shocks.
Expense-to-Income Ratio serves as a useful indicator for understanding an individual’s capacity to invest. Lower ratios imply surplus funds that can be directed toward higher-risk, long-term investments Farrell (2006), contributing to financial flexibility and resilience.
Investment Objectives and Goal Realism play a crucial role in shaping responsible investor behavior. Stanley and Danko argue that setting realistic targets proportional to income helps prevent overreaching and excessive risk-taking Stanley and Danko (1996). Unrealistic goals can encourage dangerous financial decisions, particularly among retail investors.
Emergency Funds act as a financial buffer during market downturns. Maintaining liquidity to cover at least three to six months of expenses enables investors to avoid premature liquidation of long-term assets Larimore et al. (2009). This security fosters greater risk tolerance in asset allocation.
Portfolio Diversification, as formalized in Modern Portfolio Theory, is essential in shaping both actual and perceived risk capacity. Markowitz demonstrated that spreading investments across asset classes reduces unsystematic risk while optimizing returns for a given level of volatility Markowitz (1952).
Together, these variables provide a comprehensive picture of an individual’s risk profile, integrating regulatory expectations, behavioral insights, and portfolio theory. This multi-pronged perspective guided our selection of user features in evaluating AI model performance on investment risk appetite estimation.
3 Investment Risk Scoring
We quantify the impact of these user features on investment risk by constructing a risk score from a structured scoring system that evaluates five core dimensions: Personal & Financial Stability (PFS), Investment Strategy & Objectives (ISO), Liquidity & Asset Allocation (LAA), Market & Currency Risks (MCR), and Dependency on Investments (DOI).
The Total Risk Score (TRS) is calculated using the following formula:
[…]
Based on the Raw Score, users are categorized into five risk profiles: Conservative (--14–5), Moderately (6–15), and Aggressive (16–28).
3.1 Personal & Financial Stability
This category assesses an individual’s financial stability and capacity for risk. Younger individuals (under 30) receive 2 points for their higher risk tolerance, while those over 50 generally receive 0. Users without dependents gain 2 points due to greater financial flexibility, whereas those with dependents get -1. High income (above \$100K) earns 2 points, while low income (under \$50K) results in –1. Debt-to-asset ratios below 20% indicate strong financial footing (+2), whereas ratios above 40% suggest financial vulnerability (–2). Finally, users with expense-to-income ratios below 30% suggests readiness for risk (2 points), while exceeding 50% indicates constraints (-2).
3.2 Investment Strategy & Objectives
This category evaluates the investor’s investment goals and strategy. A longer investment horizon (over 15 years) earns 2 points for higher risk tolerance, while a shorter horizon scores 0. Aggressive investment strategies (e.g., market speculation) also reflect higher risk appetite (2 points), whereas income-oriented, conservative approaches earn -1. For target amounts, staying below five times one’s income is rewarded with 2 points, while exceeding ten times leads to -2. Lastly, contributing less than 10% of income monthly suggests lower risk tolerance (2 points), whereas investing more than 30% implies higher exposure and results in -2.
3.3 Liquidity & Asset Allocation
This category evaluates the user’s ability to handle losses and manage investments. Individuals with a liquid net worth above \$500K (2 points) show strong financial stability, while net worth below \$100K (-1) signals vulnerability. Allocating more than 50% of liquid assets to investments indicates high-risk appetite (2), while allocations below 25% suggests caution (-1). Maintaining an emergency fund provides a buffer (1), while lacking one indicates reduced capacity (0).
3.4 Market & Currency Risks
This category measures exposure to market and currency volatility. Investors using only USD face lower currency risks (1 point), while those dealing with multiple foreign currencies encounter higher fluctuations (0). Extensive investment knowledge indicates higher risk tolerance (2), whereas limited experience results in a more cautious approach (-2). Lastly, diversification across asset classes mitigates risk (2), while low diversification increases vulnerability (-1).
3.5 Dependency on Investments
This final category measures financial reliance on investment returns for day-to-day expenses. Individuals who depend on investments for under 25% of their expenses gain 1 point, while those relying on 25% or more demonstrate increased vulnerability (0).
By combining scores across these five dimensions, we derive a holistic measure of each user’s investment risk appetite.
USA Australia Sweden Portugal Singapore India China Indonesia Nigeria Brazil $\Delta$
Low Tolerance Scenario (-14 to 5, Ideal = -5)
[…]
Mid Tolerance Scenario (-6 to 15, Ideal = 10)
[…]
High Tolerance Scenario (16 to 28, Ideal = 21.5)
[…]
Table 1: (Closed-weight systems): Investment risk tolerance computation. Country-specific scores are reported as mean values. The $\Delta$ column shows the mean and standard deviation of the difference between each country score and the ideal score (Low: -5, Mid: 10, High: 21.5). In the Diff column, the mean is color-coded as follows: highest in red, lowest in green, second highest in orange, and second lowest in yellow; the standard deviation is similarly color-coded.
4 Collecting Diverse User Profiles
Our approach to creating these financial profiles was guided by the need for a realistic and diverse dataset that accurately represents different investment behaviors, risk levels, and financial situations. The risk tolerance ranges from -14 to 28, covering a full spectrum of investor risk profiles, their financial situation or investment mindset suggests a strong need for stability and security, to highly aggressive and engage in high-risk, high-reward strategies.
Demographics Representation in Data.
We categorize demographic summaries into two groups based on population size and economic characteristics:
  1. 1.
    Highly Populous Countries: These include India, China, Indonesia, Nigeria, Brazil. Many individuals in these regions lack access to formal banking systems, and limited financial history often leads to lower credit scores. This can introduce biases in AI-driven risk-scoring models, as they may not accurately reflect an individual’s actual financial behavior. Additionally, economic variability and informal economies contribute to greater investment risk tolerance among individuals in these nations.
  2. 2.
    Less Populous Countries: These include Canada, Australia, Sweden, Portugal, and Singapore. Despite their smaller populations, these countries share comparable economic characteristics, including strong banking infrastructures, well-regulated credit systems, and stable financial markets. Investors in these nations typically exhibit lower risk tolerance due to greater financial security, better access to diversified investment options, and stronger social safety nets.
Genders Representation in Data.
While we acknowledge that gender is non-binary and diverse in real-world settings, the current dataset is limited to the two binary gender categories, Male and Female, based on the scope and structure of data collection.
Overall, for each of the 10 selected countries, we selected two representative names per gender. For each name, we constructed 43 summaries. Thus, in total, we obtain $10\times 2\times 2\times 43=1720$
USA Australia Sweden Portugal Singapore India China Indonesia Nigeria Brazil $\Delta$
Low Tolerance Scenario (-14 to 5, Ideal = -5)
[…]
Mid Tolerance Scenario (6 to 15, Ideal = 10)
[…]
High Tolerance Scenario (16 to 28, Ideal = 21.5)
[…]
Table 2: (Open-weight systems): Investment risk tolerance computation. Country-specific scores are reported as mean values. The $\Delta$ column shows the mean and standard deviation of the difference between each country score and the ideal score (Low: -5, Mid: 10, High: 21.5). In the Diff column, the mean is color-coded as follows: highest in red, lowest in green, second highest in orange, and second lowest in yellow; the standard deviation is similarly color-coded.
5 Experimental Setup
We evaluated a range of both proprietary and open-weight language models, selected based on their popularity, accessibility, and relevance to practical deployment in financial settings:
  • (Closed-models) We analyzed models from OpenAI (ChatGPT-4o), Google (Gemini 1.5 Pro), and Anthropic (Claude 3.7 Sonnet).
  • (Open-models) When it comes to open-weight models, we study popular choices such as Llama 3.1 (70B and 405B), DeepSeek-V3, and Mistral small (24B).
We also tried other models—Llama 3.1 (8B, 70B), Llama 3.2 (3B), and DeepSeek-R1—but skipped them due to poor adherence to instructions. For instance, Llama 3.1 (3B) often produced non-integer outputs, while Llama 3.1 (8B/70B), DeepSeek-R1, and Mistral sometimes gave inconsistent results, entered text loops, or generated poorly formatted responses.
USA Australia Sweden Portugal Singapore India China Indonesia Nigeria Brazil
GPT-4o (mini) 14.53 14.61 14.76 14.56 14.51 14.76 14.36 14.80 14.44 14.42
GPT-4o 11.70 11.58 11.78 11.48 11.66 11.58 10.94 11.96 11.63 11.47
Gemini 1.5 (Pro) 13.09 12.94 13.23 13.21 12.89 12.94 13.27 13.25 13.33 12.95
Claude 3.7 (Sonnet) 12.32 11.89 12.47 12.45 12.32 12.46 12.33 12.54 12.55 12.46
DeepSeek-V3 12.15 11.85 12.10 12.24 12.05 12.16 12.28 12.24 12.34 12.12
Llama 3.1 (405B) 12.13 11.88 11.97 12.26 12.01 11.87 12.07 12.04 12.39 12.27
Llama 3.3 (70B) 13.59 13.52 13.72 13.90 13.88 13.40 13.87 14.05 14.12 14.09
Mistral small (24B) 13.83 13.58 14.21 14.27 13.68 14.22 13.70 13.61 14.06 14.34
Table 3: Average country-wise scores for each model. Rows are color-coded to highlight the minimum (green) and maximum values (red) and their second-lowest/highest.
6 Results and Discussions
6.1 Correctness Analysis
Are Closed Models Good?
Table 1 compares each model’s deviation from the ideal risk-tolerance score in three scenarios: Low ($-5$), Mid ($10$), and High ($21.5$). We observe that GPT-4o demonstrates the smallest deviation from the ideal for both the Low (mean difference $=7.27$) and Mid ($1.84$) profiles, outperforming GPT-4o (mini), Gemini 1.5 (Pro), and Claude 3.7 (Sonnet) in those ranges. However, for the High-risk scenario, GPT-4o’s deviation ($-0.86$) is larger (i.e., farther from the ideal) than that of the other models, which lie between $+0.43$ and $-0.27$. This suggests GPT-4o is well-calibrated for low- and mid-risk profiles but less aligned with high-risk cases, suggesting the need for further refinement to capture high-risk user preferences more accurately.
Other models show distinct patterns in risk tolerance prediction. For example, GPT-4o (mini) generally exhibits larger deviations from the ideal in the low and mid scenarios, suggesting it tends to overestimate risk tolerance for lower-risk profiles. However, in the high-risk scenario its predictions are closer to the ideal, which may indicate a calibration bias that favors higher risk levels. Meanwhile, both Gemini 1.5 (Pro) and Claude 3.7 (Sonnet) deliver more moderate deviations overall. Although they do not match GPT-4o’s precision for low and mid profiles, their performance remains relatively stable across the risk spectrum, with both models showing smaller gaps in the high-risk scenario. These findings suggest that different models emphasize different dimensions of risk evaluation, implying that model selection (or even ensemble approaches) could be optimized depending on the specific risk profile.
Are Open Models Good?
Table 2 reports the performance of four open-weight systems (DeepSeek-V3, Llama 3.1 (405B), Llama 3.3 (70B), and Mistral small (24B)) across the same Low, Mid, and High risk-tolerance scenarios. We see that Llama 3.1 (405B) best matches the ideal in the Low scenario, with a mean difference of $7.00$ versus $8.10$ for DeepSeek-V3, $8.89$ for Llama 3.3, and $11.49$ for Mistral small. In the Mid scenario, DeepSeek-V3 provides the smallest gap from the target ($2.02$), followed by Llama 3.1 ($2.56$), Mistral small ($4.21$), and Llama 3.3 ($4.92$). In the High scenario, DeepSeek-V3 again shows the closest alignment to the ideal (difference $=-0.16$), with Llama 3.1 at $0.21$, Mistral small at $-0.34$ (absolute gap of $0.34$), and Llama 3.3 at $1.12$. These results indicate that while each open-weight model has its own strengths and weaknesses, DeepSeek-V3 generally performs well for mid and high risk levels, whereas Llama 3.1 (405B) excels in the low-risk setting. The standard deviations reported in the table further indicate that Mistral small and Llama 3.3 exhibit somewhat greater variability in their predictions, especially under mid and high scenarios, pointing to possible calibration gaps for more extreme risk profiles.
6.2 Consistency Analysis.
Table 1 also reports the standard deviation of each model’s predictions (relative to the ideal) across the ten countries. A lower standard deviation suggests that a model is making more consistent predictions across countries. In the Low scenario, GPT-4o (mini), GPT-4o, and Gemini 1.5 (Pro) all exhibit relatively low variability (0.25–0.26), whereas Claud 3.7 (Sonnet) is slightly higher at 0.30. In the Mid scenario, GPT-4o shows the largest cross-country spread (0.45), whereas Gemini 1.5 (Pro) has the smallest (0.13), indicating that GPT-4o’s predictions may be accurate for some countries but deviate more for others, while Gemini remains more uniform. For the High scenario, Claud 3.7 achieves the lowest standard deviation (0.17), followed by Gemini 1.5 (0.21), GPT-4o (mini) (0.24), and GPT-4o (0.31). These differences highlight each model’s varying stability across geographic contexts.
(Open Models).
In the Low scenario (Table 2), DeepSeek-V3 exhibits the smallest standard deviation (0.28), indicating relatively uniform predictions across countries, whereas Llama 3.3 (70B) has the largest spread (0.38). In the Mid scenario, Llama 3.1 (405B) is the most consistent (0.20), while Mistral small (24B) has the widest cross-country variance (0.41). Finally, for the High scenario, Llama 3.1 again shows the smallest standard deviation (0.14), whereas Llama 3.3’s predictions vary the most (0.35).
6.3 Country-Level Bias Analysis.
Table 3 indicates that no single country is universally favored or disfavored across models, though certain trends emerge. For instance, Nigeria and Indonesia often elicit higher risk-tolerance scores (e.g., for Gemini1.5, Claude .7, DeepSeek-V3, and Llama3.3), while Australia and India frequently rank near the lower end (e.g., for Claude .7, DeepSeek-V3, Mistral small). However, no country stands out as an absolute outlier across all models: China is lowest for GPT-4o (mini) and GPT-4o but mid-range elsewhere, and Australia is near the bottom for three models yet near the top for GPT-4o (mini). These mild biases likely reflect each model’s unique training or calibration, and the relatively small differences suggest no systematic disadvantage to any particular country in this dataset.
6.4 Gender Variations.
Table 4 shows gender-wise risk-tolerance scores across Low, Mid, and High scenarios with notable quantitative differences. In the Low scenario, GPT-4o (mini) assigns a male score of 8.12 in the USA—0.57 points higher than the female score of 7.55—while in Australia, the female score (7.67) surpasses the male score (7.20) by 0.47 points. Overall, GPT-4o (mini) favors males in the USA, Sweden, and Portugal, whereas GPT-4o (full) often assigns slightly higher scores to female profiles, with the exception of Indonesia. DeepSeek-V3 and Llama models generally assign higher tolerance scores to males while Mistral shows an opposite trend.
In the Mid scenario, GPT-4o (mini) shows male advantages in the USA, Singapore, China, Nigeria, and Brazil (up to +1.0) and female advantages in Australia, Sweden, Portugal, India, and Indonesia. GPT-4o (full) favors males in the USA, Australia, Portugal, and Brazil and females in Sweden, Singapore, India, and Indonesia—with differences sometimes exceeding +1.0. Gemini 1.5 (Pro) exhibits more dynamic shifts (e.g., male +1.0 in Indonesia vs. female +0.8 in Australia), while DeepSeek-V3 and the Llama models vary without a consistent gender inclination.
In the High scenario, GPT-4o (mini) tends to assign higher scores to male profiles in the USA, Australia, Portugal, Singapore, and India (differences up to +0.6) and female scores in Sweden, China, Indonesia, Nigeria, and Brazil. GPT-4o (full) shows mixed trends-favoring males in Sweden and the USA, but favoring females in Australia and India. Similarly, Gemini 1.5 (Pro), Claude 3.7 (Sonnet), and DeepSeek-V3 alternate by country, while Mistral small (24B) yields roughly +0.3–0.5 higher for males in Australia, China, and Indonesia but higher females in the USA and Portugal. Llama 3.1 (405B) generally assigns higher scores to female profiles. Overall, these differences—ranging from +0.3 to +1.0 points—underscore that gender effects vary significantly by model and region, with no uniform advantage for either gender.
7 Conclusion
This study presents a systematic evaluation of AI systems in the context of investment risk appetite assessment. We created 1,720 profiles spanning a broad spectrum of risk tolerance and varied them by country and gender. Our assessment of both closed- and open-weight models reveals notable differences in correctness, consistency, and bias across risk categories, emphasizing the need for comprehensive evaluation and alignment before deploying these systems for broader use. We provide ethics statement in Appendix.
Appendix A Appendix
Female (F) Scores Male (M) Scores
Model USA AUS SWE POR SIN IND CHN IDN NGA BRA USA AUS SWE POR SIN IND CHN IDN NGA BRA
Low Tolerance Scenario
GPT-4o (mini) 7.55 7.67 7.80 7.38 7.28 8.28 7.62 7.65 7.97 7.80 8.12 7.20 8.35 8.00 7.03 7.47 7.38 7.65 7.62 7.22
GPT-4o 2.55 2.12 2.73 2.17 2.80 2.25 1.82 2.52 2.70 1.80 2.15 2.05 2.40 2.33 2.17 1.80 1.73 2.73 1.93 2.58
Gemini 1.5 (Pro) 5.53 5.33 5.92 5.22 5.03 5.22 5.42 5.62 5.85 5.08 4.85 4.55 5.38 5.42 4.70 4.75 5.35 5.42 5.20 5.42
Claude 3.7 (Sonnet) 3.02 2.75 3.02 3.58 3.58 3.58 3.52 4.00 3.70 3.60 3.35 2.92 3.48 3.48 3.45 3.75 3.05 3.77 3.83 3.73
DeepSeek-V3 2.75 2.65 3.35 2.95 2.88 2.90 3.42 3.42 3.35 2.55 3.12 2.73 3.25 3.48 3.02 3.08 3.45 3.30 3.58 2.75
Llama 3.1 (405B) 1.95 1.45 1.27 2.25 1.45 1.93 1.93 2.05 2.10 2.62 1.95 1.60 2.00 2.35 1.98 1.88 2.25 1.90 2.52 2.58
Llama 3.3 (70B) 4.12 3.20 3.60 3.85 3.67 2.62 4.15 4.45 4.25 4.40 3.00 3.67 3.83 4.70 3.83 3.98 3.88 3.80 4.20 4.67
Mistral small (24B) 6.62 5.60 6.45 6.92 6.58 7.08 6.08 6.33 7.30 6.88 6.17 5.95 7.45 6.50 6.05 6.78 6.28 6.20 6.10 6.47
Mid Tolerance Scenario
GPT-4o (mini) 14.0 14.35 14.7 14.1 13.95 14.95 13.75 14.65 13.9 14.0 14.1 14.0 14.5 13.65 14.85 14.05 14.1 14.2 14.9 14.1
GPT-4o 11.35 11.85 11.75 11.35 12.4 12.4 10.7 12.3 12.25 11.7 12.3 12.5 11.55 11.65 11.9 11.3 10.75 12.25 12.4 12.1
Gemini 1.5 (Pro) 12.9 13.4 13.1 12.75 12.6 12.6 13.15 12.3 13.0 12.6 13.2 12.6 12.85 13.05 13.0 13.15 13.0 13.3 13.1 12.7
Claude 3.7 (Sonnet) 12.75 11.75 12.45 12.35 12.45 12.8 12.55 12.4 12.3 12.2 12.4 12.35 13.15 12.55 12.05 12.3 12.55 12.5 12.7 12.5
DeepSeek-V3 11.95 11.65 11.95 12.0 12.4 12.3 11.8 12.4 12.2 12.35 12.05 11.75 11.3 12.95 11.5 12.4 11.55 11.95 11.75 12.2
Llama 3.1 (405B) 12.55 11.85 11.9 12.8 12.25 12.05 12.4 12.2 12.55 12.55 12.65 13.25 13.4 12.75 13.0 12.2 12.65 12.9 13.15 12.1
Llama 3.3 (70B) 15.5 14.4 14.25 15.2 14.6 13.95 14.65 15.65 15.5 15.6 15.1 15.0 14.9 14.8 15.75 15.35 14.6 15.15 15.1 14.65
Mistral small (24B) 14.15 14.3 14.55 14.75 15.05 14.05 14.3 13.75 13.9 14.95 13.35 13.65 13.55 15.4 13.0 14.5 14.0 13.8 14.3 14.8
High Tolerance Scenario
GPT-4o (mini) 21.38 22.12 21.62 21.85 21.65 21.73 21.92 22.54 22.15 21.92 22.0 22.31 21.58 22.35 21.98 22.08 21.38 22.08 22.12 21.5
GPT-4o 20.77 20.92 20.92 20.77 19.88 21.15 20.35 21.04 20.46 19.62 21.12 20.04 21.31 20.62 20.81 20.62 20.31 20.92 20.04 21.04
Gemini 1.5 (Pro) 20.88 20.69 20.85 21.19 20.96 21.19 21.23 21.27 21.42 20.77 21.15 21.08 21.27 21.62 21.04 20.73 21.46 21.58 21.42 21.12
Claude 3.7 (Sonnet) 21.12 20.65 21.38 21.38 21.15 21.15 21.04 21.31 21.38 21.27 21.27 20.92 21.31 21.35 21.27 21.19 21.27 21.27 21.42 21.46
DeepSeek-V3 21.62 21.46 21.50 21.62 21.31 20.81 21.62 21.08 21.73 21.42 21.38 20.88 21.27 20.46 21.19 21.50 21.81 21.27 21.46 21.46
Llama 3.1 (405B) 21.85 21.58 21.69 21.96 21.92 21.62 21.46 21.69 22.23 21.92 21.81 21.54 21.58 21.42 21.46 21.54 21.73 21.50 21.77 21.85
Llama 3.3 (70B) 22.46 22.15 22.12 22.0 23.42 22.31 22.65 22.81 22.62 23.04 22.54 22.92 22.77 22.46 23.08 22.27 21.77 22.65 23.42 23.0
Mistral small (24B) 20.81 21.04 21.35 20.42 20.69 21.88 20.54 20.92 21.38 20.96 21.88 20.92 21.92 21.65 20.73 21.04 21.04 20.65 21.38 21.96
Table 4: Risk tolerance scores (F = female, M = male) by country, model, and scenario. Red indicates a higher predicted risk tolerance compared to its gender counterpart. Rows are grouped into Low, Mid, and High scenarios.
A.1 Ethical Considerations
We recognize that our study does not capture the full diversity of countries and gender identities, which are critical for a complete understanding of AI bias in financial risk assessment. Due to limited resources and practical constraints, our analysis focused on a select subset of countries and employed binary gender classifications. We acknowledge that these limitations may affect the generalizability of our findings. Future research should expand on this work by incorporating a broader range of demographic factors to provide a more comprehensive view of bias in AI systems.
A.2 Related Work
The integration of LLMs into financial services has opened new avenues for personalized investment advice. However, recent studies have detected potential biases in these AI-driven recommendations, raising concerns about fairness and reliability.
Bias in LLM-Generated Investment Advice.
Recent research has shown that LLMs may exhibit product biases in investment recommendations. For example, a study revealed that LLMs showed systematic preferences for specific financial products, which could subtly influence investor decisions and potentially lead to market distortions Zhi et al. (2025). Similarly studies by, Guo et al. (2024) found that LLM-generated investment advice could unintentionally elevate portfolio risks across different risk dimensions, emphasizing the importance of understanding risk biases in AI systems.
Furthermore, the inheritance of human biases by AI systems, such as home bias, can significantly affect the objectivity of financial advice Liu et al. (2024). These biases can introduce errors in asset allocation and market analysis, which can have long-term consequences for investment strategies.
Bias in AI-Driven Financial Services.
Bias in AI systems is not limited to LLMs but also extends to other AI applications in the financial sector. Studies have shown that AI systems used in areas such as mortgage underwriting and loan processing may perpetuate racial or demographic biases, resulting in discriminatory financial outcomes Raji and Buolamwini (2020). These biases, if unchecked, can detrimentally affect, the trust and reliability consumers place in financial AI models.
Frameworks for mitigating biases in AI-systems such as fairness-aware evaluations and auditing practices have been proposed by researchers Garg and Ghosh (2022). These methods focus on detecting and addressing the biases that can influence the decision-making processes in financial AI models.
Mitigation Strategies.
Efforts to mitigate biases in AI-driven financial advice include strategies such as adversarial training Madras et al. (2018), debiasing embeddings Zhao et al. (2019), and fairness-aware fine-tuningDai et al. (2023). These techniques aim to neutralize biased outcomes and improve the fairness and transparency of AI systems. Additionally, AI frameworks that offer explainability, such as Lundberg and Lee (2017) and Ribeiro et al. (2016), are also being applied to financial AI systems to make their decisions more understandable.
Our study builds upon these findings by specifically focusing on biases in LLM-generated investment advice. We aim to develop a systematic framework for evaluating these biases to eventually improve the fairness and reliability of AI-driven financial recommendations.
A.3 Country-Level Bias Analysis.
Table 3 shows the average scores (over Low, Mid, and High scenarios) for each model across ten countries. A quick inspection suggests no single country is consistently favored or disfavored by all models. However, some noteworthy patterns emerge:
  • Nigeria and Indonesia often rank near the top. For instance, Gemini 1.5 (Pro), Claude 3.7 (Sonnet), DeepSeek-V3, and Llama 3.3 (70B) each place Nigeria as either their highest or second-highest average. Meanwhile, GPT-4o (mini) and GPT-4o both reach their highest scores for Indonesia rather than Nigeria, yet Nigeria remains in the upper range for these models too when it comes to risk appetite.
  • Australia and India frequently appear near the lower end. Several models (e.g., Claude 3.7 (Sonnet), DeepSeek-V3, Mistral small (24B)) place Australia as their lowest or second-lowest country when computing risk appetite. India also appears at or near the minimum for GPT-4o, Llama 3.1 (405B), and Llama 3.3 (70B).
  • No universal outlier. Despite these recurring patterns, there is no single country that all models treat as an extreme high or low. For example, China is the lowest for GPT-4o (mini) and GPT-4o, but ranks in the middle for other systems. Similarly, Australia is near the bottom for three models yet near the top for GPT-4o (mini).
Overall, while certain countries (e.g., Nigeria, Indonesia) tend to elicit higher predicted risk-tolerance scores for multiple models, and others (e.g., Australia, India) more often appear at the lower end, these tendencies are not universal. Instead, each model exhibits its own mild biases or calibration nuances, suggesting that country-specific differences may reflect the underlying training data or optimization strategies.
A.4 Gender-Level Bias Analysis
(Low Scenario.)
  • GPT-4o (mini) tends to assign higher male scores in the USA, Sweden, and Portugal (differences of about +0.5 to +0.6), while female scores are higher in Australia, Singapore, India, China, Nigeria, and Brazil (often by +0.4 to +0.8).
  • GPT-4o similarly shows mixed results: it favors males in Portugal, Indonesia, and Brazil but females in the remaining seven countries.
  • Other models (e.g., Gemini 1.5, Claude 3.7) also exhibit alternating patterns, with some leaning toward female scores in certain countries (like Australia, Sweden) and toward male scores in others (like Portugal, Nigeria).
Mid Scenario.)
  • GPT-4o (mini) is higher for males in the USA, Singapore, China, Nigeria, and Brazil, whereas females lead in Australia, Sweden, Portugal, India, and Indonesia (differences up to +1.0).
  • GPT-4o mostly favors male scores in the USA, Australia, Portugal, China, Nigeria, and Brazil, while female scores prevail in Sweden, Singapore, India, and Indonesia.
  • The remaining systems (DeepSeek-V3, Llama variants, etc.) also produce no consistent advantage for either gender overall; rather, they alternate depending on country-specific inputs.
(High Scenario.)
  • GPT-4o (mini) shows higher male scores for the USA, Australia, Portugal, Singapore, and India (often by +0.2 to +0.6), while female scores outpace males in Sweden, China, Indonesia, Nigeria, and Brazil (differences up to +0.6).
  • GPT-4o again alternates: male scores exceed female ones in the USA, Sweden, Singapore, and Brazil, but the pattern reverses for Australia, Indonesia, and a few others.
  • Models like Gemini 1.5, Claude 3.7, and Mistral small similarly oscillate across countries, with some leaning toward higher female scores in, say, Portugal or Indonesia, and higher male scores in places like Australia or China.
Across all three scenarios, no model consistently favors one gender across all countries. Instead, the direction and magnitude of differences (often +0.3 to +0.8 points) depend on both the specific model and the user’s country profile. These findings underscore that any gender-related biases are not uniform but rather context-dependent—varying by scenario, region, and model.